# AI-Empowered Integrated Sensing and Communications

### v1

@@ III-A1 Weighted Sum Method
where $\eta$ ($0\leq\eta\leq 1$) and $\bar{\eta}=1-\eta$ are the normalizing and weighting factors assigning the priorities for sensing and communication tasks, respectively, and $f_{\rm sens}({\rm sensing\;metric})$ and $f_{\rm comm}({\rm communication\;metric})$ are some functions of sensing and communication metrics, respectively. The weights can be adjusted adaptively with respect to the requirements imposed by different scenarios. Various sensing and communication metrics can then be used.
Only some simplified forms of the problem above under some simplifying assumptions, which may not hold in practice, can be solved using optimization-based approach.
@@ III-A2 $\epsilon$-Constraint Method
In this method, one objective is optimized, while the other objectives are treated as constraints with acceptable bounds ($\epsilon$). By adjusting the $\epsilon$-values, different Pareto-optimal solutions can be generated. This method is particularly useful when there is a priority among objectives as it prioritize one objective while constraining the others.
@@
$\text{Minimize}\quad f_{i}(\cdot)$ Subject $\text{Subject to}\quad f_{j}(\cdot)$ $\leq\epsilon_{j},\quad j\in\{{\rm sens},{\rm comm}\}\backslash\{i\}$ ${g}(\cdot)$ $=0,\quad$ $\quad{h}(\cdot)$ $\leq 0,\quad$ (5)
@@
The $\epsilon$-constraint method can generate solutions on non-convex parts of the Pareto front, which the weighted sum method may miss. Also, unlike the weighted sum method, it does not require selecting weights for different objectives, thus avoiding the potential bias introduced by weight choices. This makes the $\epsilon$-constraint method more flexible and suitable for problems where it is challenging to assign relative importance to objectives.
@@ III-A3 Pareto Optimization
This approach, instead of finding a single optimal solution, aims to find the Pareto front—a set of solutions where no objective can be improved without degrading another [45, 46]. This can be achieved using techniques like evolutionary algorithms and Pareto-based methods (see, for example [47]).
@@ III-B Learning-Based Techniques
III-B1 Multi-Task Learning
Multi-task learning (MTL) is an ML paradigm where a model learns to perform multiple tasks simultaneously. Each task has its own set of outputs, and the model learns to predict those outputs jointly. The main idea is that learning shared representations across tasks can improve generalization and performance on each individual task. In contrast, single-task learning focuses on training a model for one specific task at a time, without considering interactions or dependencies with other tasks. Single-task learning is suitable when the goal is to solve a single, well-defined problem without considering the interactions or dependencies with other tasks. It is commonly used when the tasks are distinct and unrelated, or when there is insufficient data or computational resources to train a multi-task model. In multi-task learning, the goal is to train a model to perform multiple tasks simultaneously. Figure 3 illustrates the difference between these two types of learning.
@@ III-B1 Multi-Task Learning
Goal: The goal is to optimize performance on multiple tasks using shared information or features.
@@
III-B2 Multi-Objective Learning
Similar to multi-objective optimization, multi-objective learning (MOL) aims to optimize multiple, possibly conflicting, objective functions simultaneously. These objectives represent different criteria that need to be balanced within the same task. The goal is to find a balance between these objectives, often resulting in a set of solutions known as the Pareto front.
@@ III-B2 Multi-Objective Learning
where $\mathcal{L}_{i}$ is the loss for objective $i$, $\;i=1,\cdots,O$, $\alpha_{i}$ are weights, and $O$ is the number of objectives. Alternatively, other methods like the $\epsilon$-constraint method or Pareto optimization can be used to solve the multi-objective optimization problem without explicitly combining the losses into one scalar function.
@@
Goal: The goal is to optimize multiple ‘conflicting’ objectives simultaneously. The model does not simply combine the losses but tries to balance them, often leading to Pareto optimal solutions where improving one objective worsens another.
@@
In summary, multi-task learning is about performing multiple tasks simultaneously and using shared representations to improve performance. Multi-objective learning is about balancing multiple conflicting objectives during the optimization process.
@@
It is natural to treat sensing and communication as two distinct tasks, making ISAC problems inherently multi-task learning problems. However, multi-task learning problems can also be approached through multi-objective optimization [48]. In addition, in traditional multi-objective optimization, sensing and communication are considered as separate objectives, blurring the line between multi-task learning and multi-objective optimization in the context of ISAC. As such, several works, such as [49, 50, 3], have treated sensing and communication as distinct objectives.
@@ IV-A1 Waveform in Radar
  • OFDM Waveforms: Originally developed for communication systems, OFDM waveforms have been adapted for radar because of their flexibility in spectrum shaping and potential for joint radar-communication applications.
@@ IV-A2 Waveform in Communication
In communication systems, a waveform serves as the carrier of information from a transmitter to a receiver, encoding data into a format that can be efficiently transmitted over a communication channel and decoded at the receiver. Waveform design in communication involves selecting parameters such as modulation scheme, symbol rate, carrier frequency, bandwidth, and pulse shape to optimize efficiency and reliability of the transmission for the intended application and environmental conditions. OFDM has been the primary waveform used in cellular networks starting from 4G.
@@ IV-A3 Waveform in ISAC
In ISAC, a single unified waveform is used to accomplish both sensing and communication tasks. Communication waveforms can be directly employed for sensing. The multicarrier waveforms based on OFDM have shown good performance in both sensing and communication without requiring substantial alterations to existing communication system designs [55], [56]. Other waveforms such as OTFS [57] and orthogonal delay-Doppler division multiplexing [58] are shown to better suit to time-varying channels, especially in high-mobility scenarios. For example, OTFS shows greater robustness to Doppler shifts compared to OFDM in communication scenarios [59], while both waveforms exhibit comparable performance for sensing applications [60]. The difficulty with such communication-centric waveforms is that their peak-to-average power ratio (PAPR) is greater than one, indicating that the envelope of the transmitted time-domain signal does not have a constant amplitude. For example, the expected PAPR of OFDM increases with the number of subcarriers, while the expected PAPR pf OTFS increases with the number of subsymbols transmitted over slow-time in one radar frame. As a result, a non-uniform envelope requires using highly linear power amplifiers at the transmitter to avoid signal distortion that affects both radar and communication performance. It conflicts radar requirements, since high power is always beneficial for sensing. Indeed, the attenuation of a signal power for both way propagation in sensing channel is bigger than that for one way propagation in communications channel. OTFS suffers from this drawback in ISAC, just like OFDM. To deal with such difficulties, AI-based solutions have shown their importance within this approach based on the use of communication-centric waveforms [61, 62].
@@
Using the frame structure in Fig. 4, the general ISAC performance optimization problem that aims to achieve the best possibly sensing performance by using the least possible resources (number of preambles in Fig. 4) can be formulated as
@@ IV-B Beamforming Design
A review of 25 years of progress in beamforming, tracing its evolution from convex and nonconvex optimization techniques to modern machine learning approaches, with applications in radar, communications, and other domains can be found in [70]. In the following, we describe different beamforming methods used in communication, radar, and ISAC.
@@ IV-B1 Beamforming in Communication
TABLE VI: Comparison of Digital Beamforming (DBF) Techniques in Communication.
DBF Tech. Description   Advantages & Disadvantages Beamforming Formula Ref.
[…]
ZF Eliminates multi-user interference by inverting the channel matrix. + Simple implementation, effective at high SNR. Amplifies noise at low SNR; performance degrades in noisy environments. $\mathbf{W}_{\rm ZF}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H})^{-1}$ [3]
MMSE Balances interference suppression and noise amplification by utilizing a regularization factor $\alpha$ in the channel inversion. + Performs well at both low and high SNR, more robust than ZF. Higher computational complexity than ZF. $\mathbf{W}_{\rm MMSE}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H}+\alpha\mathbf{I})^{-1}$ [4]
MRT/MRC Maximizes SNR by transmitting/receiving in the direction of the strongest channel gain. + Maximizes received power, simple to implement. Does not cancel interference; not suitable for multi-user scenarios. $\mathbf{w}_{\rm MRT}=\frac{\mathbf{h}\dagger}{\|\mathbf{h}\|}$ [74]
@@
DBF in communication uses various techniques depending on system goals, complexity, and constraints. Singular value decomposition (SVD) and eigenvalue decomposition (EVD) provide optimal performance in single-user MIMO, while zero-forcing (ZF) and MMSE beamforming are widely used in multi-user MIMO to handle interference. Maximum ratio transmission (MRT) and maximum ratio combining (MRC) are simple techniques, but they lack interference mitigation, whereas Capon beamforming, also known as minimum variance distortionless response (MVDR) beamforming, is effective for adaptive spatial filtering. With emerging ML–based sensing and beamforming, future beamforming techniques are expected to become increasingly environment-aware, adaptive, and intelligent, paving the way for ISAC and 6G systems. Common DBF techniques are summarized in Table VI.
@@ IV-B3 Beamforming in ISAC
An ISAC beamformer must optimize sensing and communication accounting for trade-offs. In [83], a joint active beamforming design for base station (BS) and passive beamforming design for intelligent reflecting surfaces was proposed to enhance both communication and localization performance in the downlink scenario. Similarly, in [84], the authors studied ISAC beamforming design by maximizing the MI between the target response matrix of a point radar target and the echo signals, while ensuring the data rate requirements of the communication users. Three types of echo interference were considered: no interference, point interference, and extended interference.
@@
Interference management has also been explored in other studies. In [85], the authors investigated the precoder optimization problem to maximize radar sensing SINR, considering both signal-independent interference and signal-dependent clutter. The proposed approach ensured a minimum SINR at multiple communication users while maintaining per-antenna power constraints. Additionally, in [86], a communication-centric subcarrier allocation strategy was developed for wideband ISAC systems, where a subset of subcarriers was tailored for radar sensing while the entire bandwidth was used for communications. An efficient hybrid beamforming design was proposed on the basis of this allocation scheme.
@@
To improve spectral efficiency, researchers have explored full-duplex ISAC systems. For example, a full duplex ISAC system where the base station performs target detection while communicating with both downlink and uplink users using the same time and frequency resources was studied in [87]. An energy-efficient robust beamforming design for a full-duplex ISAC system under norm-bounded CSI errors was introduced in [88]. In [89], a new communication receiver design for a downlink multi-user communication and multi-target ISAC scenario was proposed which is capable of canceling interference caused by radar signals and maximizing the minimum weighted beampattern gain at the sensing angles of interest.
@@
Robustness under imperfect CSI was the focus of [90], where the authors studied the robust DFRC beamforming design taking into account outage SINR probability under imperfect CSI. In [91], a multi-beam design scheme was proposed where one beam is used to communicate with a single-antenna user in the downlink while one or multiple dedicated sensing beams are used to sense the unknown and random angle parameter of a target. Similarly, [92] introduced a beam pattern modulation-embedded hybrid transceiver design for mmWave ISAC systems, where the ISAC transmitter provides multi beams for single-user communication and scanning beams for sensing.
@@
In [93], transmit beamforming was studied for a multi-target, multi-user MIMO-ISAC system, where a general form and tight upper bound of sensing MI were derived, along with a multi-objective optimization framework based on user SINR and sensing MI. In [94], to mitigate interference between target echo signals while capturing near-field channel characteristics in a full-duplex ISAC system, the authors designed optimal transmit beamforming and receive beamforming, and developed an iterative beamforming algorithm to solve the resulting optimization problem. In [95], the authors exploited both the training and transmission signals for sensing, and proposed two MSE-based schemes for training and transmission design.

### v2

@@ III-A1 Weighted Sum Method
where $\eta$ ($0\leq\eta\leq 1$) and $\bar{\eta}=1-\eta$ are the normalization and weighting factors that assign priorities for the sensing and communication tasks, respectively, and $f_{\rm sens}({\rm sensing\;metric})$ and $f_{\rm comm}({\rm communication\;metric})$ are some functions of the sensing and communication metrics, respectively. The weights can be adaptively adjusted with respect to the requirements imposed by different scenarios. Various sensing and communication metrics can then be used.
Example: The weighted-sum objective approach is widely used in ISAC designs, e.g., in an ISAC-enabled base station (BS) tracking user mobility while maintaining high-rate communication [61]. Prior works have also applied this method to minimize interference and beampattern mismatch under power constraints [62], maximize combined sensing and communication mutual information with spatio-temporal power allocation [63, 64, 65], design precoders for MIMO systems to optimize achievable rates [66], and jointly tune phase-shift matrices and precoders for metrics such as SNR, SINR, rates, and CRLB [67, 68].
Remark 1.
The weighted sum fails on the non-convex Pareto front, [^2] Pareto front represents the set of solutions where no objective can be improved without worsening at least one other objective. missing some optimal points. Thus, only some simplified forms of (2) under some simplifying assumptions, which may not hold in practice, can be solved using this method. However, weighted sum is often simpler and faster to compute, especially for linear or convex problems.
@@ III-A2 $\epsilon$-Constraint Method
Introduced in [69], this method optimizes one objective while treating the remaining objectives as constraints with acceptable bounds ($\epsilon$). By adjusting the $\epsilon$-values, different Pareto-optimal solutions can be generated [58, 59]. This method is particularly useful when there is a priority among objectives, as it prioritize one objective while constraining the others.
@@
$\text{Minimize}\quad f_{i}(\cdot)$ $\text{Subject to}\quad f_{j}(\cdot)$ $\leq\epsilon_{j},\quad j\in\{{\rm sens},{\rm comm}\}\backslash\{i\}$ ${g}(\cdot)$ $=0,\quad$ $\quad{h}(\cdot)$ $\leq 0,\quad$ (3)
@@
Example: Many $\epsilon$-constraint optimization problems appear in the ISAC literature. Examples include probabilistic constellation shaping for ISAC that maximizes mutual information under a sensing constraint [70], and conversely, [71] which maximizes the sensing estimation rate under a communication constraint; SNR/CRB-constrained joint beamforming and reflection design [72]; and communication SINR-constrained CRLB minimization beamforming [73], among others.
Remark 2.
The $\epsilon$-constraint method offers an advantage over weighted-sum scalarization, as it can capture Pareto-optimal solutions in non-convex regions of the Pareto front [59]. Moreover, it does not require the selection of objective weights, reducing potential bias. However, this benefit comes at the cost of selecting and systematically varying constraint bounds ($\epsilon$-values) and potentially increased computational effort.
@@ III-A3 Pareto Optimization
This approach, instead of finding a single optimal solution, aims to find the Pareto front. This can be achieved using techniques like evolutionary algorithms and Pareto-based methods. Note that both the weighted-sum and $\epsilon$-constraint methods can generate Pareto-optimal solutions [59]. However, the weighted-sum approach can obtain only those on the convex regions of the Pareto front, whereas the $\epsilon$-constraint method is capable of identifying all Pareto-optimal solutions, including those in non-convex regions.
Example: Several ISAC studies have employed this method for optimization [74, 75, 76, 77, 78]. The motivation is to avoid reducing inherently multi-objective ISAC problems to single-objective formulations. Scalarization techniques can be problematic because the optimal solution is often subjective, being sensitive to the predefined scalar weights or constraint bounds. Also, in nonlinear and non-convex problems, such approaches may fail to capture all Pareto-optimal solutions.
@@ III-B Learning-Based Techniques
III-B1 Multi-Task Learning (MTL)
MTL is an ML paradigm where a model learns to perform multiple tasks simultaneously. Each task has its own set of outputs, and the model learns to predict those outputs jointly. The main idea is that learning shared representations across tasks can improve generalization and performance on each individual task. In contrast, single-task learning focuses on training a model for one specific task at a time, without considering interactions or dependencies with other tasks. Single-task learning is suitable when the goal is to solve a single, well-defined problem without considering the interactions or dependencies with other tasks. It is commonly used when the tasks are distinct and unrelated, or when there is insufficient data or computational resources to train a multi-task model. In multi-task learning, the goal is to train a model to perform multiple tasks simultaneously. Figure 2 illustrates the difference between these two types of learning.
@@ III-B1 Multi-Task Learning (MTL)
The goal is to optimize performance on multiple tasks using shared information or features.
@@
III-B2 Multi-Objective Learning (MOL)
Similar to multi-objective optimization, MOL aims to optimize multiple, possibly conflicting, objective functions simultaneously. These objectives represent different criteria that need to be balanced within the same task. The goal is to find a balance between these objectives, often resulting in a set of solutions known as the Pareto front.
@@ III-B2 Multi-Objective Learning (MOL)
where $\mathcal{L}_{i}$ is the loss for objective $i$, $i=1,\dots,O$, $\alpha_{i}$ are weights, and $O$ is the number of objectives. Alternatively, other methods like the $\epsilon$-constraint method or Pareto optimization can be used to solve the multi-objective optimization problem without explicitly combining the losses into one scalar function.
@@
The goal is to optimize multiple ‘conflicting’ objectives simultaneously. The model does not simply combine the losses but tries to balance them, often leading to Pareto optimal solutions where improving one objective worsens another.
@@
In summary, multi-task learning is about performing multiple tasks simultaneously and using shared representations to improve performance. Multi-objective learning is about balancing multiple conflicting objectives during the optimization process. The boundaries between multi-task learning and multi-objective learning in ISAC are scenario-dependent, as explained below.
@@
The distinction between multi-task learning and multi-objective learning in ISAC is often scenario-dependent and conceptually subtle. When sensing and communication are treated asdistinct tasks producing separate outputs (e.g., channel estimation and target detection), the formulation follows the MTL paradigm. Conversely, when the focus lies on balancing conflicting objectives within a single system (e.g., optimizing rate–sensing trade-offs), the problem aligns more closely with MOL. In practice, the boundary between the two is blurred: ISAC systems can be modeled as MTL problems while still being optimized via multi-objective formulations [79]. Indeed, several recent works treat sensing and communication as distinct objectives within a multi-objective optimization framework [80, 81, 3]. Therefore, the distinction ultimately depends on whether the emphasis is placed on the MTL or the MOL.
@@ IV-A1 Waveform in Radar
  • OFDM Waveforms: Originally developed for communication systems, OFDM waveforms have been adapted for radar because of their flexibility in spectrum shaping and potential for joint radar-communication applications [86].
@@ IV-A2 Waveform in Communication
In communication systems, a waveform serves as the carrier of information from a transmitter to a receiver, encoding data into a format that can be efficiently transmitted over a communication channel and decoded at the receiver. Waveform design in communication involves selecting parameters such as modulation scheme, symbol rate, carrier frequency, bandwidth, and pulse shape to optimize efficiency and reliability of the transmission for the intended application and environmental conditions.
OFDM has been the primary waveform in modern wireless systems, including WiFi and cellular networks since 4G. One main drawback of OFDM is the high peak-to-average power ratio (PAPR), the ratio between the peak power and the average power of a signal. A waveform with high PAPR requires highly linear power amplifiers to avoid signal distortion. Such amplifiers are typically less power-efficient, leading to increased hardware costs and higher energy consumption—particularly challenging for battery-powered devices.
In full-duplex systems, the BS transmits and receives simultaneously over the same frequency band, introducing additional waveform design constraints due to self-interference, spectral coexistence, and hardware limitations. High-PAPR waveforms impose stringent linearity and dynamic-range requirements on full-duplex transmitters. When self-interference suppression is limited, the residual self-interference after analog cancellation is often dominated by power amplifier-induced nonlinear distortion, which scales with PAPR. Consequently, waveforms with lower PAPR, or employing effective PAPR reduction techniques, can indirectly enhance self-interference suppression performance by reducing nonlinear residual components. Conversely, higher self-interference suppression capability relaxes power amplifier back-off constraints and improves the robustness of high-PAPR waveforms such as OFDM and OTFS [87, 88, 89].
@@ IV-A3 Waveform in ISAC
In ISAC, a single unified waveform is used to accomplish both sensing and communication tasks. Communication waveforms can be directly employed for sensing. The multicarrier waveforms based on OFDM have shown good performance in both sensing and communication without requiring substantial alterations to existing communication system designs [90, 52].
In OFDM-based systems like 5G, unused subcarriers or sync blocks can be repurposed as radar subcarriers to enhance sensing performance, though this may reduce SNR for communication users due to power reallocation to radar subcarriers. It can also increase PAPR due to the use of more subcarriers. Optimizing these radar subcarriers can minimize estimation error (e.g., via CRLB) and reduce PAPR, improving overall system efficiency [91]. In [92], an iterative algorithm based on the majorization-minimization technique is proposed to minimize PAPR under the constraint that the integrated periodic autocorrelation sidelobe level is zero.
Other waveforms such as OTFS [93] and orthogonal delay-Doppler division multiplexing [94] are shown to better suit to time-varying channels, especially in high-mobility scenarios. For example, OTFS shows greater robustness to Doppler shifts compared to OFDM in communication scenarios [95], while both waveforms exhibit comparable performance for sensing applications [96]. The difficulty with such communication-centric waveforms is that their PAPR is greater than one, indicating that the envelope of the transmitted time-domain signal does not have a constant amplitude. For example, the expected PAPR of OFDM increases with the number of subcarriers, while the expected PAPR pf OTFS increases with the number of subsymbols transmitted over slow-time in one radar frame. As a result, a non-uniform envelope requires using highly linear power amplifiers at the transmitter to avoid signal distortion that affects both radar and communication performance. It conflicts radar requirements, since high power is always beneficial for sensing. Indeed, the attenuation of a signal power for both way propagation in sensing channel is bigger than that for one way propagation in communications channel. OTFS suffers from this drawback in ISAC, just like OFDM. To deal with such difficulties, AI-based solutions have shown their importance within this approach based on the use of communication-centric waveforms [97, 98].
@@
Full duplex communication can increase spectral efficiency. It, however, introduces additional waveform design constraints due to self-interference, spectral coexistence, and hardware limitations, all of which can significantly impact the PAPR of transmitted signals, particularly in ISAC systems [91, 86]. Particularly, for self-interference suppression, waveforms must be designed to minimize leakage into the receiver path [105]. This may require power shaping or signal clipping, which can increase PAPR due to waveform distortion.
@@
Using the frame structure in Fig. 3, the general ISAC optimization problem that aims to achieve the best possibly sensing performance by using the least possible resources (number of preambles in Fig. 3) can be formulated as
@@ IV-B Beamforming Design
A review of 25 years of beamforming progress—from convex and non-convex optimization to modern machine learning approaches, with applications in radar, communications, and beyond—can be found in[107]. In the following, we describe different beamforming methods used in communication, radar, and ISAC.
@@ IV-B1 Beamforming in Communication
TABLE V: Comparison of Digital Beamforming (DBF) Techniques in Communication
Technique Description Advantages & Disadvantages Beamforming Formula Ref.
SVD Uses SVD to decompose the correlated channel $\mathbf{H}$ into a set of orthogonal channels. + Eliminates interference between antennas; optimal for single-user MIMO. Requires perfect CSI; computationally expensive. $\text{Channel:}\;\mathbf{H}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{H},$ $\mathbf{W}_{\text{SVD}}=\mathbf{V}$ [108, 109]
EVD Uses eigenvalue decomposition on the signal covariance matrix $\mathbf{Q}$ to extract dominant spatial directions for beamforming. + Effective for spatial filtering and directional beamforming. Requires statistical knowledge of the channel; high computational cost. $\text{Covariance:}\;\mathbf{Q}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{H},$ $\mathbf{W}_{\text{EVD}}=\mathbf{V}$ [110]
ZF Eliminates multi-user interference by inverting the channel matrix. + Simple implementation, effective at high SNR. Amplifies noise at low SNR; performance degrades in noisy environments. $\mathbf{W}_{\text{ZF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H})^{-1}$ [3]
MMSE Balances interference suppression and noise amplification by utilizing a regularization factor $\alpha$ in the channel inversion. + Performs well at both low and high SNR, more robust than ZF. Higher computational complexity than ZF. $\mathbf{W}_{\text{MMSE}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H}+\alpha\mathbf{I})^{-1}$ [4]
MRT/MRC Maximizes SNR by transmitting/receiving in the direction of the strongest channel gain. + Maximizes received power, simple to implement. Does not cancel interference; not suitable for multi-user scenarios. $\mathbf{w}_{\text{MRT}}=\frac{\mathbf{h}^{\dagger}}{\|\mathbf{h}\|}$ [111]
@@
DBF in communication uses various techniques depending on system goals, complexity, and constraints. Singular value decomposition (SVD) and eigenvalue decomposition (EVD) provide optimal performance in single-user MIMO, while zero-forcing (ZF) and MMSE beamforming are widely used in multi-user MIMO to handle interference. Maximum ratio transmission (MRT) and maximum ratio combining (MRC) are simple techniques, but they lack interference mitigation, whereas Capon beamforming, also known as minimum variance distortionless response (MVDR) beamforming, is effective for adaptive spatial filtering. With emerging ML–based sensing and beamforming, future beamforming techniques are expected to become increasingly environment-aware, adaptive, and intelligent, paving the way for ISAC and 6G systems. Common DBF techniques are summarized in Table V.
@@ IV-B3 Beamforming in ISAC
An ISAC beamformer must optimize sensing and communication accounting for trade-offs. In [120], a joint active beamforming design for BS and passive beamforming design for intelligent reflecting surfaces was proposed to enhance both communication and localization performance in the downlink scenario. Similarly, in [121], the authors studied ISAC beamforming design by maximizing the MI between the target response matrix of a point radar target and the echo signals, while ensuring the data rate requirements of the communication users. Three types of echo interference were considered: no interference, point interference, and extended interference.
@@
Interference management has also been explored in other studies. In [122], the authors investigated the precoder optimization problem to maximize radar sensing signal-to-interference-plus-noise ratio (SINR), considering both signal-independent interference and signal-dependent clutter. The proposed approach ensured a minimum SINR at multiple communication users while maintaining per-antenna power constraints. Additionally, in [123], a communication-centric subcarrier allocation strategy was developed for wideband ISAC systems, where a subset of subcarriers was tailored for radar sensing while the entire bandwidth was used for communications. An efficient hybrid beamforming design was proposed on the basis of this allocation scheme.
@@
To improve spectral efficiency in ISAC, researchers have explored ISAC systems with full-duplex communication. [^3] Note that sensing is inherently full-duplex as a sensing BS transmits and receives radar signals simultaneously. However, communication can operate in downlink, uplink, or both directions. The latter corresponds to full-duplex communication, while the former two represent half-duplex communication. For example, beamforming design for an ISAC system, where the BS performs target detection while simultaneously communicating with two different groups of users in the uplink and downlink, is studied in [124]. An energy-efficient robust beamforming design for ISAC systems with full-duplex communication under norm-bounded CSI errors was introduced in [125].
@@
Robustness under imperfect CSI was studied in [126], which proposed a robust beamforming for DFRC based on outage SINR probability. In [127], a multi-beam design scheme was proposed where one beam is used to communicate with a single-antenna user in the downlink while one or multiple dedicated sensing beams are used to sense the unknown and random angle parameter of a target. Similarly, [128] introduced a beam pattern modulation-embedded hybrid transceiver design for mmWave ISAC systems, where the ISAC transmitter provides multi beams for single-user communication and scanning beams for sensing.
@@
In [129], transmit beamforming was studied for a multi-target, multi-user MIMO-ISAC system, where a general form and tight upper bound of sensing MI were derived, along with a multi-objective optimization framework based on user SINR and sensing MI. In [130], an iterative beamforming algorithm is developed to mitigateinterference between target echo signals while capturing near-field channel characteristics in an ISAC system. In [131], the authors exploited both the training and transmission signals for sensing, and proposed two MSE-based schemes for training and transmission design. In [132] holographic-surface–based ultra-massive MIMO and its application to ISAC is investigated.